# appAppendix References

### v6

@@ Abstract
Using FOMC votes and meeting transcripts from 1976–2018, we develop a deep learning model based on self-attention mechanism to quantify “hidden dissent” among members. Although explicit dissent is rare, we find that members often have reservations with the policy decision, and hidden dissent is mostly driven by current or predicted macroeconomic data. Additionally, hidden dissent strongly correlates with data from the Summary of Economic Projections and a measure of monetary policy sub-optimality, suggesting it reflects both divergent preferences and differing economic outlooks among members. Finally, financial markets show an immediate response to the hidden dissent disclosed through meeting minutes.
@@ 1 Introduction
Statements, minutes, speeches, and other texts released by the Federal Reserve have been scrutinized extensively using techniques ranging from simple word clouds to advanced machine learning models. However, the transcripts of Federal Open Market Committee (FOMC) meetings—word-for-word records of member statements—have been less studied. This is partly due to the five-year delay in their release, but also because of the difficulty in analyzing the vast volume of text produced by over a hundred past FOMC members.
@@
To validate the relevance of our measure, we regress it on two sets of explanatory variables at both individual and meeting levels. The first set includes current or projected targets under the dual mandate: inflation and unemployment rates. The second set comprises member characteristics like age, gender, political ideology, and education. Our results show that hidden dissent is strongly correlated with economic conditions, increasing when inflation rises or unemployment falls. This correlation suggests that members have more reservations when there seems to be a need for “cooling down” the economy, and members are more in unison when the economy is in a downturn. Among characteristics, we find that only education education background has consistent explanatory power.
@@
Finally, we investigate how financial markets respond to hidden dissent by focusing on meeting minutes instead of transcripts. Unlike transcripts, which are subject to a five-year embargo, the minutes are released three weeks after each meeting, offering us a sufficient and timely indication of hidden dissent and allowing us to better capture potential market reactions. Our analysis shows that the hidden dissent revealed in the minutes provides an important supplementary channel of information, complementing the sentiment conveyed through other FOMC communication channels.
@@ 1.1 Related papers
Our paper differs by using FOMC transcripts, which provide a more detailed record of members’ views. We construct a measure of hidden dissent at the meeting level rather than at the individual level, allowing members’ policy stances to vary over time rather than categorically labeling them as hawks or doves. Our findings indicate that hidden dissent within a meeting is closely related to current economic conditions, with little influence from personal characteristics.
In addition, we also apply our deep learning model to FOMC meeting minutes to derive a more timely assessment of hidden dissent. This approach takes advantage of our model’s ability to capture nuanced dissent, offering a consistent, data-driven index that can be utilized soon after meetings, rather than waiting for the five-year lagged release of transcripts.
@@ 3 Natural Language Processing
With the continued evolution of deep learning, a multitude of NLP tasks are seeing improving outcomes (Kamath et al., 2019; Lauriola et al., 2022). In this section, we present the architecture of our customized deep learning model, which is built on the self-attention mechanism. We implement the model on FOMC transcripts and members’ speeches. For background information on the application of deep learning models in NLP, please refer to Appendix A. A more detailed discussion can be found in Dell (2024).
@@ 3.1 A customized deep learning model
We use the Sentence-BERT (SBERT) model developed by Reimers and Gurevych (2019) to do sentence embedding. Diverging from the conventional word embedding, SBERT adds a pooling operation to the output of the original BERT model to produce a sentence embedding, and it can significantly reduce computational overhead. Theoretically, the “sentence” fed into the SBERT model can be a real sentence, a paragraph, or even a full article. In this paper, we apply the SBERT at the sentence level for two reasons: (1) we want to strike a balance between manageable computational overhead and maximal information retention; (2) given transcripts are aggregated at the individual level, neighboring sentences within one person’s transcript may arise from disparate conversations during the meeting, potentially lacking strong semantic cohesion. After obtaining the sentence embedding for the transcript via SBERT, we feed these embeddings into our customized deep learning model. [^4] For each transcript associated with an FOMC member in every meeting, SBERT produces an embedding tensor of shape (number of sentences, 768), where 768 represents the embedding dimension. To ensure consistency in tensor size across the dataset, we apply padding to each embedding tensor to achieve a standardized size of (256, 768). We deliberately selected the size of 256 for the standardized tensor to optimize information density inside the tensor, ensuring it is comprehensive yet not excessively sparse. These standardized embeddings are subsequently fed into our customized deep learning model.
Drawing inspiration from the versatility of the self-attention mechanism, this study introduces a novel deep learning approach that merges the capabilities of the multi-head self-attention (MHSA) module with multi-layer perceptron (MLP) layers to forecast the voting scores of members in FOMC meetings. Historical analysis of FOMC meetings reveals a consistent pattern where chairs vote affirmatively. Leveraging this insight, our training methodology employs the chair’s transcript as a reference point, focusing on the difference between the transcripts of members and their chair to predict members’ voting outcomes. Initially, sentence embeddings from both the chair’s and the member’s transcripts are processed through separate dense layers to highlight crucial attributes. This is followed by the application of a normalization layer to enhance model efficiency and a dropout layer to counteract overfitting. The incorporation of MHSA modules further optimizes the feature extraction phase, while the addition of a residual layer aims to bolster model performance and counter potential degradation issues. The model ultimately predicts the voting outcomes, categorized as YES (0) or NO (1), by feeding the output differences derived from the chair’s and member’s MHSA modules into a classification layer. The architecture of the complete model is depicted in Figure 3.
The next step is to conduct hyperparameter fine-tuning to identify the most effective configuration for our model. The selected hyperparameters for adjustment include: (1) the count of MHSA modules within the sections dedicated to the chair’s and member’s transcripts, (2) the number of attention heads within each MHSA module in both sections, (3) the dropout rate and (4) the initial learning rate. The fine-tuning process utilizes the optuna package, with the specific search domain for each hyperparameter described in Table 4.
In the FOMC voting record, the number of NO votes is significantly smaller than that of YES votes, resulting in an imbalanced dataset. To address the issues of biased predictions and misleading accuracy rates, we over-sample the under-represented category (i.e., transcripts associated with NO votes) in both the training set and test set, ensuring we have an equal number of observations from both categories. Additionally, to further enhance the model performance and expedite convergence, a learning rate scheduler is incorporated into the model training process. We employ a step-based decay strategy, reducing the initial learning rate after every ten training steps. [^5] We have also experimented with other schedulers, such as using different step thresholds or incorporating restarts into the schedulers to mitigate potential issues with local minima. However, our findings indicate that the current configuration yields the best model performance. Furthermore, to mitigate overfitting and reduce computation time, an early stopping mechanism is activated, using the loss value on the test set as the criterion for termination.
Finally, the model with the best performance is configured with the following hyperparameters: three MHSA modules for both chair and member section, each comprising eight heads; a dropout rate of 0.735; and an initial learning rate of $4.0\times 10^{-5}$. This configuration achieves an 85.88% accuracy rate on the training set (80% of the total observations) and a 84.20% accuracy rate on the test set (20% of the total observations). [^6] Technically, oversampling the test set is unnecessary; however, given the highly skewed class distribution, a balanced test set may facilitate the model training process. Our method achieves an accuracy rate of 78.41% on the original test set.
In this binary classification models (see Figure 3), a round function is employed right after the sigmoid function to convert the probability range ($[0,1]$) into a binary outcome (0 or 1). By removing the final round function in the aforementioned deep learning model, we get the probability of a member’s transcript being predicted as 1 (i.e., vote for no, different from the chair’s vote). Hence, the prediction score extracted from the sigmoid function can be interpreted as how close the member agrees with the chair, and we call it the hidden dissent score from now on. We denote $hd_{ij}$ as FOMC member $j$’s hidden dissent score in meeting $i$. If $hd_{ij}$ is close to 0, it means that member $j$ is unlikely to vote NO and leans to agree with the chair in FOMC meeting $i$. Conversely, if $hd_{ij}$ is close to 1, this member has a high chance to disagree and vote NO. It is important to note that achieving a high accuracy rate is not the sole objective in this study. Given the severe imbalance in the original dataset, a naïve approach that predicts every observation as 0 would yield an accuracy above 90%. However, such predictions fail to reveal any meaningful insights into hidden dissent. In section 4, we demonstrate that $hd_{ij}$ provides unique information not captured by other FOMC communication channels.
@@ 3.2 Hidden dissent inside FOMC meeting transcripts
With the hidden dissent scores ($hd_{ij}$) from our customized model, we can now study what characteristics contribute to FOMC members’ voting behaviors. Moreover, we are able to describe the hidden dissents that cannot be fully captured by voting records ($v_{ij}=0$ or $1$).
@@ 4.1 What drives the hidden dissent inside FOMC meetings?
$y_{ij}=\beta_{0}+\bm{\beta_{1}}X_{i}^{MacroVariables}+\bm{\beta_{2}}X_{ij}^{PersonalCharacters}+\varepsilon_{ij},$
where $y_{ij}=hd_{ij},v_{ij}$. Considering the unbalanced nature of the panel data and the conditions where $hd_{ij}\in(0,1)$ and $v_{ij}\in\{0,1\}$, we employ mixed-effects beta panel regression for $hd_{ij}$ and mixed-effects logistic panel regression for $v_{ij}$, respectively. [^8] For these and all the following regressions, using OLS yields similar results and does not change our conclusions. Both regressions are clustered at the individual level and the findings are presented in Table 6, showing that certain macroeconomic variables and personal characteristics can help explain the hidden dissent. In particular, as inflation is trending up, members on average have more reservations about the FOMC decision. This observation echos the conclusion from (Ball, 1992) that high inflation may raise inflation uncertainty since when inflation is high, the Federal Reserve is facing a dilemma: curbing inflation or a potential recession. Consistent with Bordo and Istrefi (2023), where the member went to school matters. Regarding other personal characteristics, the findings remain inconclusive, leading us to consider these factors as control variables. For a comprehensive presentation of the results, please refer to Appendix D. Furthermore, if we regress on voting results, $v_{ij}$, instead, while results for education mostly remain, macroeconomic variables do not matter anymore.
Next, we conduct a similar regression at the meeting level to examine the relationship between the hidden dissent level $HD_{i}$ for meeting $i$, macroeconomic variables, and the composition of the FOMC committee’s characteristics in meetings. The committee’s characteristics at the meeting level are summarized as follows:
@@ 4.2 Is hidden dissent about what will be or what should be?
In our analysis, we categorize SEP variables into two groups: projections of the federal funds rate, which represent what the member considers as the “appropriate monetary policy”, and projections of other economic indicators, which indicate the members’ expectations for the economic outlook. To quantify disagreement embedded within these projections, we calculate the absolute sum of deviations from the median projection for each category, following the approach used by Foerster and Martinez (2023). Due to the limited sample size and high correlations among the projections, we consolidate the SEP policy disagreement measures across different horizons (current year, next year, the following year, and long term) into up to two principal component, which explains $83.36\%$ of the variations. For the economic outlook-related disagreement measures (same horizons, except no long run values for inflation measures), we use up to three principal components to capture the corresponding SEP disagreement levels, and they can explain $84.16\%$ of the total variance.
We then regress the hidden dissent measure derived from meeting transcripts against each category of SEP disagreements, and the results are shown in Table 8. First, disagreement regarding the path of the federal funds rate clearly predicts the hidden dissent identified in our transcripts, evidenced by an $R^{2}$ of $0.1531$ (column (1)). Second, disagreement about economic forecasts has about the same explanatory power: while the $R^{2}$ is lower in the longer sample that begins in 2007 in column (3), it is higher using the same sample as the federal funds rate in column (4). [^11] If the absolute sum of deviations from the mean projection is utilized to compute SEP disagreements, we still see similar results. See Table B.3. These findings suggest that our hidden dissent measure is equally associated with the heterogeneity in members’ policy preferences and their views on the economic outlooks.
SEP policy disagreement is clearly influenced by disagreement about the economy outlook, and it is possible that policy disagreement matters only because of the other disagreement. Indeed, the first component of SEP policy disagreement ($\text{Policy}_{PC1}$) and the first component of SEP economic outlook disagreement ($\text{Economy}_{PC1}$) are highly correlated, with a coefficient of $0.7323$. We would like to know the relationship between our hidden dissent measure and one category of SEP disagreement after accounting for the influence of the other.
@@
To achieve this, we apply the Double Machine Learning (DML) method proposed by Chernozhukov et al. (2017). DML combines machine learning techniques with econometric models to control for confounding variables, and it aims to isolate the causal effect of a treatment variable ($T$) on the variable of interest ($Y$). The DML process involves two stages: first, machine learning algorithms are used to predict $Y$ and $T$ separately based on confounder variables $X$, and the algorithm helps us to capture the complex relationships between $Y$ or $T$ and $X$. Next, the residuals from these predictions are regressed against each other to estimate the causal effect. In our setting, each SEP disagreement takes the role of $T$ or $X$ alternately, with our transcript hidden dissent measure ($HD_{i}$) being $Y$ [^12] SEP policy disagreement is characterized by four distinct measures, while SEP economic outlook disagreement encompasses fourteen measures. We treat the first principal component of one SEP disagreement as $T$, and utilize all measures from the other SEP disagreement as $X$. The Lasso model ($\lambda=1$) is employed in the first stage, followed by linear regression in the second stage. To mitigate overfitting and reduce estimation bias, cross-fitting is used in our model.
@@
The second-stage results are reported in columns (5) and (6) of Table 8. The significant and positive coefficient for $\text{Residualized Policy}_{PC1}$ implies that the disagreement on monetary policy significantly contributes to hidden dissent inside FOMC meetings, even after accounting for disagreement on the economy outlook. This suggests that policy disagreement encompasses additional factors related to our measure. One possibility is members having different loss functions, but it is also possible that members differ on variables not covered in the SEP. In contrast, $\text{Residualized Economy}_{PC1}$ is not statistically significant, suggesting that disagreement about the economic outlook does not have an independent influence on our measure.

### v7

# Agree to Disagree: Measuring Hidden Dissent in FOMC Meetings

@@ Abstract
Using FOMC transcripts and customized deep learning models, we quantify “hidden dissent”, or disagreement in the FOMC that is unobserved in formal votes. We find hidden dissent to be prevalent and systematically driven by macroeconomic conditions like inflation and unemployment. It strongly correlates with divergent member projections (SEP) and measures of policy sub-optimality, reflecting heterogeneity among members in policy preferences. Furthermore, we show that the financial markets respond to the hidden dissent implied in FOMC minutes.
@@ 1 Introduction
Statements, minutes, speeches, and other texts released by the Federal Reserve have been scrutinized extensively using techniques ranging from simple word clouds to advanced machine learning models. However, the transcripts of Federal Open Market Committee (FOMC) meetings—word-for-word records of members’ views—have been less studied. This is partly due to the five-year delay in their release, but also because of the difficulty in analyzing the vast volume of text produced by over a hundred past FOMC members.
@@
To validate the relevance of our measure, we regress it on two sets of explanatory variables at both individual and meeting levels. The first set includes current or projected targets under the dual mandate: inflation and unemployment rates. The second set comprises member characteristics like age, gender, political ideology, and education. Our results show that hidden dissent is strongly correlated with economic conditions, increasing when inflation rises or unemployment falls. This correlation suggests that members have more reservations when there seems to be a need for “cooling down” the economy, and members are more in unison when the economy is in a downturn. Among characteristics, we find that only education background has consistent explanatory power.
@@
Finally, we investigate the real-world relevance of this hidden dissent by examining its impact on financial markets. While previous research has focused heavily on the sentiment conveyed in central bank communications (e.g., overall hawkishness or dovishness), we hypothesize that the degree of internal hidden disagreement represents a distinct and previously under-explored information channel. Using FOMC minutes for timeliness, we demonstrate that our hidden dissent measure provides information beyond simple sentiment, triggering significant and unique reactions in stock and bond markets.
@@ 1.1 Related papers
Our paper moves beyond traditional static classifications by using FOMC transcripts, which provide a more detailed record of members’ views. By applying our deep learning model to the full content of transcripts, we construct a hidden dissent score at the individual-meeting level rather than the individual level, allowing members’ policy preferences to fluctuate over time rather than categorically labeling them as hawks or doves. This method provides a more flexible and granular view compared to assigning fixed labels, capturing the nuances of policy deliberation. We also show that hidden dissent is closely related to current economic conditions, with little influence from personal characteristics.
In addition, our work connects to the literature examining explicit voting dissent as an information channel in monetary policy. Studies like Riboni and Ruge-Murcia (2014) demonstrate that dissenting votes can predict future policy actions, while Madeira et al. (2023) link dissent frequency to specific economic conditions, such as supply shocks that create policy trade-offs. However, a key limitation of focusing solely on recorded votes is their rarity. In contrast, our hidden dissent measure proves to be more prevalent and, as we demonstrate, shows a closer connection with the macroeconomic situation and tangible influences on financial markets, complementing the insights gained from studying explicit dissent.
Furthermore, we also apply our deep learning model to FOMC meeting minutes to derive a more timely assessment of hidden dissent. This approach takes advantage of our model’s ability to capture nuanced dissent, offering a consistent, data-driven index that can be utilized soon after meetings, rather than waiting for the five-year lagged release of transcripts.
@@ 3 Natural Language Processing
With the continued evolution of deep learning, a multitude of NLP tasks are seeing improving outcomes (Kamath et al., 2019; Lauriola et al., 2022). In this section, we present the architecture of our customized deep learning model, which is built on the self-attention mechanism. We implement the model on FOMC communication materials. For background information on the application of deep learning models in NLP, please refer to Appendix A. A more detailed discussion can be found in Dell (2024).
@@ 3.1 A customized deep learning model
Our primary goal is to quantify the extent of disagreement expressed by FOMC members during meetings. To achieve this, we develop a deep learning model designed to analyze the textual content of members’ statements. Since computers process numerical data rather than raw text, the first step involves converting the language used by members into a format the model can understand. We employ a state-of-the-art technique based on Sentence-BERT (Reimers and Gurevych, 2019) to transform each sentence within a member’s transcript into a numerical representation, known as a sentence embedding. This process captures the semantic meaning of the sentences. For technical consistency, we ensure all these numerical representations have a uniform size through the padding process. [^4] Specifically, using Sentence-BERT, each sentence is converted into a 768-dimension vector. We then group these sentence vectors for each member’s transcript in a meeting and pad them to a standard size of (256, 768) before feeding them into our model. This standardization aids the model training process.
The core of our approach lies in comparing each member’s statements against those of the FOMC chair. Given the Chair’s consistent record of voting with the majority to implement the final policy decision, we use their transcript as a proxy for the meeting’s ultimate policy stance or consensus view. Our customized model architecture, inspired by the self-attention mechanism prevalent in modern NLP (Vaswani et al., 2017), is specifically designed for this comparison. Intuitively, the self-attention mechanism allows the model to weigh the importance of different parts of the text, focusing on the segments most relevant for gauging agreement or disagreement. The model processes the numerical representations of both the member’s and the chair’s transcripts, identifies the key differences in their expressed views, and uses these differences to predict the member’s likely voting outcome (YES=0 or NO=1). Figure 3 provides a schematic overview of this architecture.
Training such a model requires careful consideration of the data and methodology. We optimize the model’s configuration by fine-tuning its hyperparameters, such as the complexity of its internal layers and the learning rate, using the optuna framework developed by Akiba et al. (2019) to find the best settings (details in Table 4). A key challenge is the inherent imbalance in the voting data: dissenting (NO) votes are rare compared to assenting (YES) votes. [^5] After filtering out observations that contain fewer than five sentences, the final dataset contains 3,523 observations. Among these, only 244 (6.9%) are linked with NO votes, compared to 3,279 with YES votes. This severe imbalance poses challenges not only for current model training, but also for further distinguishing the direction of dissent (e.g., dovish vs. hawkish). With so few NO votes, further subdividing them into directional categories would leave too few observations per class for reliable prediction. Accordingly, our analysis is constrained to measuring the magnitude of dissent. Nevertheless, we show that our proposed measure of hidden dissent is a potent, and previously unmeasured, channel of information that provides statistically significant and economically meaningful insights. To prevent the model from simply learning to always predict YES, we employ an over-sampling technique, effectively showing the model an equal number of YES and NO vote examples during training. Furthermore, we incorporate standard deep learning practices like learning rate scheduling (gradually adjusting the learning rate during training) and early stopping (halting training when performance on unseen data stops improving) to enhance training efficiency and, crucially, to prevent overfitting—ensuring the model generalizes well to new, unseen transcripts rather than just memorizing the training data.
The optimally configured model achieves high accuracy in predicting actual votes on both the training data (85.88%) and unseen test data (84.20%). While the model is trained to predict a binary outcome (0 or 1), the underlying prediction is a probability score between 0 and 1, indicating the likelihood of a member dissenting (voting NO). We leverage this probability directly as our measure of hidden dissent. We denote this score as $hd_{ij}$ for member $j$ in meeting $i$. A score close to 0 signifies strong agreement with the chair’s position, while a score close to 1 indicates substantial disagreement, reflecting a high probability of the model predicting a NO vote. Summary statistics for these hidden dissent measures are reported in Table 5.This continuous measure allows us to capture nuances in members’ positions that a simple binary vote cannot. Furthermore, in section 4, we demonstrate that $hd_{ij}$ provides unique information not captured by other FOMC communication channels.
@@ 3.2 Hidden dissent inside FOMC meeting transcripts
With the hidden dissent scores ($hd_{ij}$) from our customized model, we can now study what characteristics contribute to FOMC members’ voting behaviors. Moreover, we are able to describe the hidden dissent that cannot be fully captured by voting records ($v_{ij}=0$ or $1$).
@@
3.3 Hidden dissent: hawkish or dovish?
Sentiment analysis that classifies policy preferences along a hawkish-dovish spectrum is a widely adopted and valuable method for studying the FOMC (Bordo and Istrefi, 2023; Malmendier et al., 2021; Istrefi, 2019). We propose our hidden dissent score as a complement to this approach, designed to capture a different aspect of policy deliberation. While sentiment indices effectively measure the directional leaning of preferences, we find that disagreements are often multi-dimensional and cannot be fully captured by a binary classification. Our measure therefore quantifies the intensity of these nuanced disagreements.
The limitations of a simple hawk-dove classification can be illustrated by the voting record of President Jerry Jordan. In the October 6, 1992 meeting, President Jordan dissented in favor of a more accommodative policy, a seemingly “dovish” action. However, his reasoning was not a simple call for more loosening but a critique of the Fed’s lack of credibility:
“I ask business people why they don’t believe us when we say we’re going to zero inflation…They just sort of shrug and say they don’t believe it…I think the issue we have to attack is the credibility of our commitment; we as a Committee need to convey in the clearest way we possibly can that we are going toward zero inflation. That’s what we should tell them about how to gauge our policy.”
Less than two year later, in the March 22, 1994 meeting, President Jordan dissented again, this time in favor of a more aggressive tightening (a 50 basis point hike instead of the Committee’s 25). This appears to be a classic “hawkish” dissent. His reasoning was again rooted in establishing a systematic policy to bolster credibility, arguing that the Committee was not moving decisively enough toward a neutral stance:
“And it’s a useful idea to think that there is such a thing as a neutral policy stance. I feel very strongly that we are nowhere near a neutral stance and that we ought to be aggressive in moving toward it.”
President Jordan was not simply a “dove” in 1992 and a “hawk” in 1994. Rather, he was a consistent advocate for a credible, systematic policy framework. He dissented when he believed the Committee’s actions—whether by being too timid in tightening or too discretionary in easing—failed to achieve that primary objective. Our model captures this consistency, assigning high hidden dissent scores to President Jordan in both instances (0.82 in 1992 and 0.80 in 1994), reflecting his significant deviation from the Committee’s discretionary approach, regardless of the policy direction.
Besides “No” votes that are more than hawkish or dovish, we also have “No” votes that do not have a directional leaning. According to the dataset compiled by Thornton and Wheelock (2014), out of a total of 503 dissenting votes have been recorded between March 19, 1936 and September 17, 2025, 269 are categorized as being for a tighter policy and 172 for an easier policy. The remaining 62 dissents—over 12% of the total—could not be classified in either category. [^6] Restricting the data to our 1976–2018 sample period, 30 of the 281 recorded dissents are classified as neither “tighter” nor “easier.” One example of such a dissent is that of President Charles Plosser in the September 17, 2014, meeting. He objected explicitly to the wording and communication strategy of the policy statement, not its immediate policy direction:
“At the last meeting I dissented over forward guidance language in alternative B…I have to renew my objection to the language, which continues, in my mind, to ignore the significant progress the economy has made toward our goals. It is not just the last number that matters. It is the accumulated progress we have made over the past year. And both the statement language and the forward guidance, in particular, fail to acknowledge that.”
Ultimately, our hidden dissent score complements existing hawk-dove sentiment indices by capturing these nuanced, multi-dimensional disagreements. The analysis in subsection 4.3 will further show that this measure provides additional information not captured by traditional sentiment indices.
@@ 4.1 What drives the hidden dissent inside FOMC meetings?
$y_{ij}=\beta_{0}+\bm{\beta_{1}}X_{i}^{Macro}+\bm{\beta_{2}}X_{ij}^{Char}+\varepsilon_{ij},$ (1)
where $y_{ij}=hd_{ij},v_{ij}$. Considering the unbalanced nature of the panel data and the conditions where $hd_{ij}\in(0,1)$ and $v_{ij}\in\{0,1\}$, we employ mixed-effects beta panel regression for $hd_{ij}$ and mixed-effects logistic panel regression for $v_{ij}$, respectively. [^8] For these and all the following regressions, using OLS yields similar results and does not change our conclusions. Both regressions are clustered at the individual level and the findings are presented in Table 6, showing that certain macroeconomic variables and personal characteristics can help explain the hidden dissent. In particular, as inflation is trending up, members on average have more reservations about the FOMC decision. This observation resonates with the conclusion from (Ball, 1992) that high inflation may raise inflation uncertainty since when inflation is high, the Federal Reserve is facing a dilemma: curbing inflation or a potential recession. Consistent with Bordo and Istrefi (2023), where the member went to school matters. Regarding other personal characteristics, the findings remain inconclusive, leading us to consider these factors as control variables. For a comprehensive presentation of the results, please refer to Appendix D. Furthermore, if we regress on voting results, $v_{ij}$, instead, while results for education mostly remain, macroeconomic variables do not matter anymore.
Our analysis thus far shows that, while hidden dissent is strongly related to the inflation mandate, its connection to unemployment is surprisingly modest. This is somewhat puzzling, given that unemployment is the other key mandate of the Federal Reserve. One potential explanation, advanced by Malmendier et al. (2021), is that members’ personal experiences generate heterogeneity in their sensitivity to macroeconomic variables. [^9] We thank an anonymous referee for the suggestion to explore this channel. To test this hypothesis, we conduct another analysis with interaction terms between macroeconomic variables and members’ formative-year experiences, estimated using a fixed-effects model:
$y_{ij}=\beta_{0}+\bm{\beta_{1}}X_{i}^{Macro}+\bm{\beta_{2}}X_{i}^{Macro}\cdot X_{j}^{Exp}+\tau_{j}+\varepsilon_{ij},$ (2)
where $X_{j}^{Exp}$ represents one of three events experienced by FOMC members in their early ages: the Great Depression, World War II, or the Great Inflation. $\tau_{j}$ is the fixed effect for member $j$.
The results, presented in Table 7, strongly support this hypothesis. The modest aggregate effect of unemployment masks significant heterogeneity driven by members’ early-age experiences. Specifically, during periods of high unemployment, members who, in their early age, lived through the Great Depression or WWII (two adjacent events) are significantly more likely to support the consensus accommodative policy. In contrast, members whose views were shaped by the Great Inflation exhibit significantly greater conservatism and are more likely to express dissent from policies of trading off unemployment for inflation.
This finding not only provides further support for the mechanism in Malmendier et al. (2021) but also demonstrates the value of our hidden dissent measure. When we repeat the analysis using official votes ($v_{ij}$) as the dependent variable (see Table B.1), these nuanced effects disappear entirely, which demonstrates again that the official voting record is too coarse to detect these subtle yet economically important differences in policy preferences.
Next, we conduct a similar regression as equation (1) at the meeting level to examine the relationship between the hidden dissent level $HD_{i}$ for meeting $i$, macroeconomic variables, and the composition of the FOMC committee’s characteristics in meetings. The committee’s characteristics at the meeting level are summarized as follows:
@@ 4.2 Is hidden dissent about what will be or what should be?
In our analysis, we categorize SEP variables into two groups: projections of the federal funds rate, which represent what the member considers as the “appropriate monetary policy”, and projections of other economic indicators, which indicate the members’ expectations for the economic outlook. To quantify disagreement embedded within these projections, we calculate the average absolute deviation from the median projection for each category, following the approach used by Foerster and Martinez (2023). Due to the limited sample size and high correlations among the projections, we consolidate the SEP policy disagreement measures across different horizons (current year, next year, the following year, and long term) into up to two principal component, which explains $78.61\%$ of the variations. For the economic outlook-related disagreement measures (same horizons, except no long run values for inflation measures), we use up to three principal components to capture $82.76\%$ of the total variance.
We begin by regressing our hidden dissent measure on each category of SEP disagreement separately, with the results presented in Table 9. A clear pattern emerges: disagreement over the appropriate path of monetary policy is a strong and statistically significant predictor of hidden dissent, alone explaining over $12\%$ of its variation (Column 1). In contrast, disagreement over the economic outlook is neither statistically nor economically significant, and its explanatory power is negligible in the longer sample (Column 3). This initial evidence suggests that our hidden dissent measure primarily captures heterogeneity in members’ policy preferences rather than differing views of the economic outlooks.
A potential challenge to this interpretation, however, is that the two forms of disagreement are themselves moderately correlated. The correlation of $\text{Policy}_{PC1}$ and $\text{Economy}_{PC1}$ is 0.5640-0.5640. A member’s view on the appropriate policy path is naturally informed by their economic outlook. To isolate the independent contribution of each channel and test the robustness of our initial finding, we must therefore disentangle these effects.
@@
To achieve this, we apply the Double Machine Learning (DML) method proposed by Chernozhukov et al. (2017). DML combines machine learning techniques with econometric models to control for confounding variables, and it aims to isolate the causal effect of a treatment variable ($T$) on the variable of interest ($Y$). The DML process involves two stages: first, machine learning algorithms are used to predict $Y$ and $T$ separately based on confounder variables $X$, and the algorithm helps us to capture the complex relationships between $Y$ or $T$ and $X$. Next, the residuals from these predictions are regressed against each other to estimate the causal effect. In our setting, each SEP disagreement takes the role of $T$ or $X$ alternately, with our transcript hidden dissent measure ($HD_{i}$) being YY. [^12] SEP policy disagreement is characterized by four distinct measures, while SEP economic outlook disagreement encompasses fourteen measures. We treat the first principal component of one SEP disagreement as $T$, and utilize all measures from the other SEP disagreement as $X$. The Lasso model is employed in the first stage, followed by linear regression in the second stage. Hyperparameter tuning is conducted to select the optimal model parameters. To mitigate overfitting and reduce estimation bias, cross-fitting is used in our model.
@@
The second-stage results are reported in columns (5) and (6) of Table 9. The significant coefficient for $\text{Residualized Policy}_{PC1}$ implies that the disagreement on monetary policy significantly contributes to hidden dissent inside FOMC meetings, even after accounting for disagreement on the economy outlook. This suggests that policy disagreement encompasses additional factors related to our measure. One possibility is members having different loss functions, but it is also possible that members differ on variables not covered in the SEP. In contrast, $\text{Residualized Economy}_{PC1}$ is not statistically significant, suggesting that disagreement about the economic outlook does not have an independent influence on our measure.